### v1

# How Does Quantization Affect Multilingual LLMs?

Abstract
Quantization techniques are widely used to improve inference speed and deployment of large language models. While a wide body of work examines the impact of quantized LLMs on English tasks, none have examined the effect of quantization across languages. We conduct a thorough analysis of quantized multilingual LLMs, focusing on their performance across languages and at varying scales. We use automatic benchmarks, LLM-as-a-Judge methods, and human evaluation, finding that (1) harmful effects of quantization are apparent in human evaluation, and automatic metrics severely underestimate the detriment: a 1.7% average drop in Japanese across automatic tasks corresponds to a 16.0% drop reported by human evaluators on realistic prompts; (2) languages are disparately affected by quantization, with non-Latin script languages impacted worst; and (3) challenging tasks such as mathematical reasoning degrade fastest. As the ability to serve low-compute models is critical for wide global adoption of NLP technologies, our results urge consideration of multilingual performance as a key evaluation criterion for efficient models.
How Does Quantization Affect Multilingual LLMs?
Kelly Marchisio1, Saurabh Dash2, Hongyu Chen1, Dennis Aumiller1, Ahmet Üstün2, Sara Hooker2, Sebastian Ruder1 1Cohere 2Cohere For AI Correspondence: kelly@cohere.com
1 Introduction
Multilingual large language models (LLMs) have the power to bring modern language technology to the world, but only if they are cheap and reliable. Known as the low-resource double bind, underserved languages and severe compute constraints often geographically co-occur (Ahia et al., 2021), meaning that for wide adoption, multilingual LLMs must be highly-performant and lightweight.
With the shift towards large models, quantization is a widely adopted technique to reduce cost, improve inference speed, and enable wider deployment of LLMs. Work on quantization, however, is by-and-large evaluated in English only (e.g. Xiao et al., 2023; Ahmadian et al., 2024; Frantar et al., 2022). No works to our knowledge have characterized the impact of quantization on the multilingual generation capabilities expected from modern LLMs. Ubiquitous use of compression techniques in the real world drives urgency to the question how are multilingual models impacted?
Our question is timely, given recent work showing that compression techniques such as quantization and sparsity amplify disparate treatment of long-tail features, which may have implications for under-represented languages in multilingual LLMs (Hooker et al., 2019, 2020; Ahia et al., 2021; Ogueji et al., 2022). Indeed, many model designs choices implicitly overfit to a handful of resource rich languages: from tokenizer choice, to weighting of training data, and to widely-used quantization techniques. Focusing on a small subset of high-resource languages in design degrades model performance for overlooked languages (Schwartz et al., 2022; Kotek et al., 2023; Khandelwal et al., 2023; Vashishtha et al., 2023; Khondaker et al., 2023; Pozzobon et al., 2024), introduces security vulnerabilities (Yong et al., 2023; Nasr et al., 2023; Li et al., 2023a; Lukas et al., 2023; Deng et al., 2023), and unfairly passes high costs to non-English users faced with high latency (Held et al., 2023; Durmus et al., 2023; Nicholas and Bhatia, 2023; Ojo et al., 2023; Ahia et al., 2023).
We analyze four state-of-the-art multilingual LLMs across 3 different sizes ranging from 8 to 103 billion parameters and covering up to 23 languages, under various quantization techniques. Critically, it is vital that we move beyond automatic evaluation and gather real human feedback on performance cost. We thus perform multilingual human evaluation on challenging real-world prompts in addition to LLM-as-a-Judge and evaluation on standard automatic benchmarks such as multilingual MMLU Hendrycks et al. (2020), MGSM Shi et al. (2023), and FLORES-200 (Costa-jussà et al., 2022a). Across experimental set-ups we find that:
  1. 1.
    Automatic metrics severely underestimate damage from quantization. Automatic evaluations estimate performance deterioration relative to FP16 across tasks at $-0.3\%$ (French) and $-1.7\%$ (Japanese) vs $-16.6\%$ and $-16.0\%$ reported by human evaluators.
  2. 2.
    Quantization affects languages differently. Non-Latin script languages are more greatly harmed on average. Across tasks, Latin-script languages scored $-0.7\%$ relative to FP16 for a 103B parameter model while non-Latin scripts scored $-1.9\%$. For a smaller 8-billion parameter model, scores were $-3.0\%$ vs. $-3.7\%$.
  3. 3.
    Challenging tasks degrade fastest. Mathematical reasoning ($-13.1\%$), performance on real-world challenging prompts judged by humans ($-10.5\%$), and LLM-as-a-Judge ($-25.9\%$) results are severely degraded.
  4. 4.
    Occasionally, quantization brings benefits. Similar to Ahia et al. (2021) and Ogueji et al. (2022) on sparsity, we find that quantization benefits model performance in some cases: e.g., an average 1.3% boost across tasks for a 35B model quantized with W8A8.
As the first to broadly study the impact of quantization on multilingual LLMs, our work is part of a wider body of literature that considers the impact of model design choices on downstream performance. Our results urge attention to multilingual performance at all stages of system design.
2 Background
Quantization compresses the weights and potentially activations of a neural network to lower-bit representations. Compression can be done by training the model at lower precision, known as Quantization Aware Training (QAT), or performed on the final model weights, known as Post Training Quantization (PTQ). Given the difficulties in training LLMs especially at precision lower than 16-bits floating point, PTQ methods which perform the quantization single-shot without needing gradient updates are highly desirable. Training is completed at higher precision, then weights/activations are quantized without further training. In this work, we focus on post-training quantization because of its simplicity and applicability at scale. PTQ of LLMs can be further categorized into:
Weight-Only Quantization
Weight matrices are quantized offline and the compressed matrices are loaded from memory during inference. Quantized weight matrices have a smaller memory footprint compared to FP16 (2×2$\times$2 × smaller for 8-bit and almost 4×4$\times$4 × smaller for 4-bit), enabling inference with less compute. In memory-bound scenarios, it also enables faster inference due to fewer bytes transferred from GPU memory to the compute units.
For a weight matrix $\mathbf{W}\in\mathbb{R}^{d_{in}\times d_{out}}$ and input $\mathbf{X}\in\mathbb{R}^{seq\times d_{in}}$, if only a single scaling factor is used for naive quantization (per-tensor), then the quantized weights are given by:
$\mathbf{W}_{Q}=\Delta\cdot\left\lfloor\frac{\mathbf{W}}{\Delta}\right\rceil,\quad\Delta=\frac{\max(|\mathbf{W}|)}{2^{N-1}}$ (1)
[…]
A single scaling factor might not be enough if the distribution of parameters in the weight matrix has high variance; thus one could increase the granularity of quantization by using a scale for each output dimension (per-column), i.e., $\Delta\in\mathbb{R}^{d_{out}}$. However, when $N$ is aggressively lowered to 4 bits or lower, even per-column granularity might be insufficient to cover the range of values in a column. The granularity can be further increased by using a shared scale for a subset of input dimensions called groups ($g$), thus the scale $\Delta\in\mathbb{R}^{\frac{d_{in}}{g}\times d_{out}}$. A commonly used group size is 128.
Equation 1 gives the simplest way to quantize the weights. For $N\leq 4$ bits, using more advanced Weight-Only Quantization methods like GPTQ (Frantar et al., 2022) or AWQ (Lin et al., 2024) leads to better downstream performance.
Weight-and-Activation Quantization
As the name suggests, Weight-and-Activation Quantization quantizes the model activations alongside the weights. Unlike Weight-Only Quantization where weights can be quantized offline, quantization of activations happens at runtime. One could compute the quantization scales for various activations by using a small slice of training or validation data (static scaling) but this method typically has large degradation (Xiao et al., 2023). For minimal degradation, it is preferred to calculate the quantization scaling factor dynamically (dynamic scaling) for each input on-the-fly. While quantizing activations is more difficult, reducing the precision of the activations alongside the weights enables the usage of specialized low-precision matrix multiplication hardware in modern GPUs leading to up to 2×2$\times$2 × improvement in throughput. For a weight matrix $\mathbf{W}\in\mathbb{R}^{d_{in}\times d_{out}}$ and input $\mathbf{X}\in\mathbb{R}^{seq\times d_{in}}$, naive Weight-and-Activation Quantization with per-token input granularity and per-column weight granularity generates output $\mathbf{Y}\in\mathbb{R}^{seq\times d_{out}}$ by:
${\mathbf{W}_{Q}}_{:,j}=\left\lfloor\frac{\mathbf{W}_{:,j}}{\Delta^{W}_{:,j}}\right\rceil,\Delta^{W}_{:,j}=\frac{\max(|\mathbf{W}_{:,j}|)}{2^{N-1}}$ (2)
${\mathbf{X}_{Q}}_{i,:}=\left\lfloor\frac{\mathbf{X}_{i,:}}{\Delta^{X}_{i,:}}\right\rceil,\Delta^{X}_{i,:}=\frac{\max(|\mathbf{X}_{i,:}|)}{2^{N-1}}$ (3)
$Y=\Delta^{X}\odot(\mathbf{X}_{Q}\mathbf{W}_{Q})\odot\Delta^{W}$ (4)
where, $\Delta^{W}\in\mathbb{R}^{d_{out}}$, $\Delta^{X}\in\mathbb{R}^{seq}$ and $\odot$ denotes element-wise multiplication by broadcasting the elements to match the shape of the operands.
3 Experiment Set-up
Models
We evaluate Command R+ [^1] https://docs.cohere.com/docs/command-r-plus, Command R [^2] https://docs.cohere.com/docs/command-r, and Aya 23 models (Aryabumi et al., 2024) as representatives of state-of-the-art multilingual LLMs. Command models are 103 and 35 billion parameters, and Aya 23 models are 35 and 8 billion parameters. We quantize the models using the weights available on HuggingFace.
Quantization
For Command R and R+, we evaluate both weight-only quantization at 8-bit (W8 with per-column scaling) and 4-bit (W4-g with group-wise scaling using GPTQ (Frantar et al., 2022)), as well as weight-and-activation quantization at 8-bit (W8A8 with per-column scaling for weights and per-token scaling for activations).
Ahmadian et al. (2024) show that if the model is trained with the right hyper-parameters, naive Weight-and-Activation Quantization has minimal degradation. Otherwise, one may leverage SmoothQuant (Xiao et al., 2023) to smoothen the distribution of activations making them more amenable to quantization. We therefore also explore W8A8-SmoothQuant (a W8A8 variant with SmoothQuant) for Command R+ as well as a 4-bit weight-only quantized variant with column-wise scaling (W4) to understand the impact of scaling granularity at extremely low-bit precision. Following (Frantar et al., 2022; Xiao et al., 2023), we use 128 English samples for calibration for SmoothQuant and GPTQ.
For Aya 23 8B and 35B, we use bitsandbytes [^3] https://github.com/TimDettmers/bitsandbytes to obtain 8-bit and 4-bit quantized models. Bitsandbytes uses LLM.int8() (Dettmers et al., 2022)—similar to W8A8 described above except that it performs certain computations in FP16. Bitsandbytes 4-bit uses the NF4 datatype (Dettmers et al., 2023) to perform Quantile Quantization which limits degradation at the expense of inference speedups.
3.1 Automatic Evaluation
We evaluate in 10 primary languages: Arabic, French, German, English, Spanish, Italian, Portuguese, Korean, Japanese, and Chinese. Quantized models are compared to the original FP16 versions, and we primarily report results as relative degradation compared to this FP16 baseline:
%$\%\Delta=\frac{\text{score}_{\text{quantized}}-\text{score}_{\text{FP16}}}{\text{score}_{\text{FP16}}}*100$ (5)
Raw numeric results are in the Appendix. Results are averaged over 5 runs.444k=0, p=0.75, temp=0.3, except mMMLU, which, as a QA eval, is run deterministically with temp=0.
Multilingual MMLU (mMMLU)
This multi-domain question answering task consists of 14,000+ multiple-choice questions. We translate MMLU (Hendrycks et al., 2020) to 9 languages with Google Translate and refer to this version as mMMLU. We measure accuracy in a 5-shot setting. An example is in Table A1.
MGSM (Shi et al., 2023)
MGSM is a generative mathematics evaluation set manually translated from GSM8K (Cobbe et al., 2021). Of our target languages, it is available for German, Spanish, French, Japanese, and Chinese. We report accuracy over the 250-item test set for each language.
FLORES-200 (Costa-jussà et al., 2022b)
This well-known multi-way parallel test set evaluates translation capabilities. We translate into and out of English, and report SacreBLEU Post (2018).
Language Confusion (Marchisio et al., 2024)
These test sets assess a model’s ability to respond in a user’s desired language. In the monolingual setting, prompts are in language $l$ and the model must respond in language $l$. For instance, a user prompts in Arabic, so implicitly desires an Arabic response. In the cross-lingual variant, a prompt is provided in English but the user requests output in a different language $l^{\prime}$. [^5] An example from the Okapi subsection of the evaluation is: “Reply in Spanish. Explain a common misconception about your topic. Topic: Using AI to Augment Human Capabilities” fastText Joulin et al. (2016) language identification is run over the output. We report line-level pass rate (LPR), i.e., the percentage of responses for which all lines in the response are in the user’s desired language.
Aya Evaluation
Aya 23 models are evaluated using an extended version of the Aya evaluation setup (Aryabumi et al., 2024) using the unseen discriminative tasks (XWinograd (Muennighoff et al., 2023), XCOPA (Ponti et al., 2020), XStoryCloze (Lin et al., 2022)), mMMLU (Okapi; Dac Lai et al., 2023), MGSM, and Belebele (Bandarkar et al., 2023) from eval-harness (Gao et al., 2023). [^6] We follow the setup used by Üstün et al. (2024): each evaluation is run once, and for FLORES, no sampling is used and metric is spBLEU. We evaluate models on languages included in the covered 23 languages, except for the unseen tasks where we use all available languages.777mMMLU: ar, de, es, fr, hi, id, it, nl, pt, ro, ru, uk, vi, zh. MGSM: de, es, fr, ja, ru, zh. Belebele: {mMMLU} + cs, fa, el, ja, ko, pl, tr. FLORES: {Belebele} + he. Aya evaluations allow us to add: Czech, Greek, Hebrew, Hindi, Indonesian, Dutch, Persian, Polish, Romanian, Russian, Turkish, Ukrainian, Vietnamese.
3.2 Human Evaluation
We run human evaluation in Spanish, French, Korean, and Japanese.
Internal Evaluation Suite
150 diverse prompts designed to be more complex than public evaluation benchmarks. As such, we expect greater degradation with increased quantization given the difficulty of the samples. Prompts for all four languages are translated by humans from an English seed prompt, ensuring that respective language-specific subsets share the same prompts.
Aya Dolly-200  (Singh et al., 2024)
We use multilingual data from the Aya Evaluation Suite to assess open-ended generation capabilities. For Korean and Japanese, we use prompts from the Aya Dolly-200 test set (dolly-machine-translated), which are automatically translated from English Dolly-15k (Conover et al., 2023) then human-curated to avoid references requiring specific cultural or geographic knowledge. For French and Spanish, we use dolly-human-edited, a human post-edited version of dolly-machine-translated. For each language, we evaluate using the first 150 prompts.
Annotator Statistics
Annotations and translations were completed by native-level speakers of the respective languages, each of whom is also fluent in English. Annotators were paid by the hour, with compensation above the federal minimum wage of the country of employment.
Annotation Interface
We use a pairwise evaluation setup. Annotators see a prompt and two (shuffled) completions of the FP16 model and a quantized variant. They rate each response on a 5-point Likert scale, then express a preference between the two model outputs (tie, weak preference, strong preference). We encourage annotators to avoid tied rankings. Win rates are based on the ranking preferences alone.
3.3 LLM/RM-as-a-Judge
Because human evaluation is costly and time-intensive, it is common to use an “LLM-as-a-Judge” to rate model completions (e.g. Li et al., 2023b; Zheng et al., 2023). Reward models (RMs) can also simulate human preference. An RM scores multiple completions given the same prompt, and the prompt-completion pair with the higher score is deemed preferred. We call this RM-as-a-Judge.
We assess quantized model outputs using LLM- and RM-as-a-Judge. In the former, an LLM selects a preferred response from a $<$instruction, modelA_completion, modelB_completion$>$ tuple (see Table A2). Following Üstün et al. (2024); Aryabumi et al. (2024) we use GPT-4888gpt-4-turbo (gpt-4-1106-preview): https://platform.openai.com/docs/models/gpt-4-turbo-and-gpt-4 as an LLM proxy judge. To minimize bias, we randomize the order of model outputs. For RM-as-a-Judge, a multilingual reward model scores $<$prompt, completion$>$ pairs for each model output, over which we calculate win-rate. We report win-rates of quantized models versus the FP16 baseline.
We assess the outputs of quantized models over the Internal Evaluation Suite and Aya Dolly-200 described in Section 3.2. We use the same prompt and completion pairs as in human evaluation, which provides the ability to relate LLM/RM-as-a-Judge performance with human evaluation.
4 Results
Avg. FLORES Language Confusion
[…]
FP16 - 66.7 - 70.6 - 37.7 - 39.6 - 99.2 - 91.5 -
W8 -0.2% 66.7 0.0% 69.9 -1.0% 37.7 0.0% 39.6 0.0% 99.2 0.0% 91.2 -0.3%
W8A8-sq -0.5% 66.3 -0.5% 69.5 -1.6% 37.8 0.2% 39.1 -1.3% 99.2 0.0% 91.5 0.1%
W8A8 -0.8% 65.6 -1.7% 69.8 -1.1% 37.7 0.0% 39.1 -1.2% 99.4 0.2% 90.4 -1.2%
W4-g -0.9% 65.7 -1.4% 68.6 -2.9% 37.8 0.4% 39.4 -0.5% 99.2 0.0% 90.5 -1.1%
103B W4 -2.5% 63.8 -4.3% 64.4 -8.8% 37.1 -1.6% 39.0 -1.6% 99.3 0.1% 92.8 1.4%
FP16 - 59.4 - 49.8 - 32.4 - 35.5 - 98.7 - 66.5 -
W8 -0.2% 59.3 -0.1% 49.4 -0.7% 32.3 -0.2% 35.4 -0.2% 98.8 0.1% 66.3 -0.2%
W8A8 0.2% 59.3 -0.2% 47.1 -5.5% 32.9 1.6% 35.8 0.9% 99.0 0.3% 68.9 3.7%
35B W4-g -2.8% 58.2 -2.0% 43.3 -13.1% 31.7 -1.9% 35.3 -0.7% 98.3 -0.4% 67.1 1.0%
Table 1: Per-dataset average performance across non-English languages for 103B and 35B Command models at varying levels of quantization. %$\Delta$ the relative performance vs. FP16 [ex., for MGSM at W4-g on the 35B: $\frac{43.3-49.8}{49.8}*100=-13.1\%$.] Languages: ar, de, es, fr, it, ja, ko, pt, zh; except MGSM: de, es, fr, ja, zh. Any discrepancy is due to rounding: raw scores and %$\Delta$ were calculated at full precision.
Avg. FLORES Unseen
[…]
FP16 - 58.2 - 51.2 - 37.8 - 42.9 - 77.6 - 70.8 -
W8 0.1% 57.9 -0.5% 52.1 1.8% 37.9 0.3% 43.0 0.1% 77.1 -0.6% 70.6 -0.2%
Aya 35B W4 -2.9% 56.6 -2.7% 48.1 -6.0% 37.2 -1.4% 42.4 -1.2% 73.0 -5.9% 70.5 -0.3%
FP16 - 48.2 - 34.7 - 34.8 - 39.5 - 64.8 - 67.6 -
W8 0.3% 47.8 -0.9% 35.4 2.1% 34.8 0.2% 39.7 0.5% 64.6 -0.3% 67.6 0.1%
Aya 8B W4 -3.7% 46.7 -3.2% 32.1 -7.5% 34.1 -1.8% 39.1 -1.0% 59.3 -8.5% 67.5 -0.2%
Table 2: Per-dataset average performance across non-English languages for 35B and 8B Aya 23 models at varying levels of quantization. %$\Delta$ is relative performance vs. FP16. We follow the evaluation setup of Aryabumi et al. (2024) and evaluate on languages in the 23 languages list. On “Unseen Tasks” (XWinograd, XCOPA, XStoryCloze), we use all the available languages. See Section 3.1 for details and language list.
To clearly see the many-faceted impact of quantization, we discuss our results by quantization level (§4.1), by task (§4.2), by language (§4.3), by model size (§4.4), and by quantization strategy (§4.5). We then report LLM-as-a-Judge and RM-as-a-Judge (§4.6) and human evaluation results (§4.7).
4.1 By Quantization Level
How do different levels of quantization affect downstream performance?
Command R and R+
In Table 1, we aggregate results of each metric for each level of quantization. We average scores across languages, then calculate the relative percentage drop from FP16.999Ex. For 103B W4-g MGSM, scores were: {de: 71.2, es: 75.7, fr: 69.0, ja: 58.0, zh: 68.9}, thus the average score was 68.6—a 2.9% drop from FP16 ($\frac{68.6-70.6}{70.6}=-0.029$). We discuss results of W8, W8A8, and W4-g quantization, which are variants available for both Command model sizes. Most results follow intuition: greater quantization leads to larger performance degradation: $-0.2\%$ for W8, $-0.8\%$ for W8A8, and $-0.9\%$ overall for the 103B model. An exception is W8A8 for the 35B, which experiences a slight boost overall due to higher performance on translation and language confusion evaluations.
Aya 23 Models
Table 2 shows the aggregated results for Aya 23 models on the extended Aya evaluations at W8, and W4 quantization. We find a similar trend with Command models where W4 often leads to a larger drop compared to W8, consistent across tasks and languages. W8, however, does not substantially drop performance in any task.
4.2 By Task
Are tasks differently affected by quantization?
Results here reference Tables 1 and 2, with full raw and relative results in Appendix A.2. Mathematical reasoning as measured by MGSM is strikingly affected by quantization. Relative performance of the 35B W4-g model is a dismal $-13.1\%$ on average, with as poor as $-17.3\%$ in Chinese. MGSM and Belebele are most greatly degraded for Aya 23 models with W4 quantization, dropping $7.5\%$ and $8.5\%$ on the 8B model. mMMLU is the next most greatly degraded task.
ar de es fr it ja ko pt zh Avg Ltn/IE $\neg$
W8 0.0% 0.1% 0.0% 0.0% 0.0% 0.1% 0.0% -0.4% -0.2% -0.1% -0.1% -0.1%
W8A8-sq -0.6% 0.2% -0.3% 0.1% -0.6% -0.3% -0.1% -0.7% -0.8% -0.3% -0.3% -0.4%
W8A8 -1.3% -0.9% -0.5% -0.5% -0.8% -0.3% -1.3% -0.8% -0.9% -0.8% -0.7% -0.9%
W4-g -0.8% -0.2% -0.4% 0.1% -0.4% -0.4% -0.6% -1.2% -0.9% -0.5% -0.4% -0.7%
103B W4 -1.0% -0.6% 0.1% -0.8% -1.2% -1.4% -2.9% -0.8% -2.3% -1.2% -0.7% -1.9%
W8 0.3% -0.5% -0.1% -0.2% -0.4% 0.3% -0.1% 0.1% -0.3% -0.1% -0.2% 0.0%
W8A8 2.0% 2.5% 0.7% 1.0% 1.2% 1.1% 0.9% 1.4% 1.0% 1.3% 1.3% 1.3%
35B W4-g -1.1% -1.1% 0.1% -0.3% -0.1% -2.3% -1.4% -0.6% -1.3% -0.9% -0.4% -1.5%
Table 3: Per-language relative performance (%$\Delta$) vs. FP16, averaged over mMMLU, FLORES, and Language Confusion tasks. Ltn/IE are Latin-script/Indo-European languages: de, es, fr, it, pt. $\neg$ are the rest: ar, ja, ko, zh.
de es fr ja zh Avg Ltn/IE $\neg$
W8 0.1% -0.1% -0.3% -0.4% -0.2% -0.2% -0.1% -0.3%
W8A8-sq 0.4% -0.9% -0.1% -0.3% -1.2% -0.4% -0.2% -0.8%
W8A8 -0.4% -1.0% -0.6% -0.1% -1.3% -0.7% -0.6% -0.7%
W4-g -0.5% -0.5% -0.3% -1.7% -1.1% -0.8% -0.4% -1.4%
103B W4 -2.3% -1.1% -1.7% -3.0% -3.5% -2.3% -1.7% -3.3%
W8 -0.6% -0.3% -0.1% -0.4% 0.0% -0.2% -0.3% -0.2%
W8A8 1.3% -0.6% 0.3% -0.3% 0.0% 0.1% 0.3% -0.2%
35B W4-g -3.7% -1.8% -1.7% -3.8% -4.0% -3.0% -2.4% -3.9%
Table 4: Per-language relative performance (%$\Delta$) vs. FP16, averaged over MGSM, mMMLU, FLORES, and Language Confusion tasks. Ltn/IE are Latin-script/Indo-European: de, es, fr. $\neg$ are the rest: ja, zh.
[…]
Curiously, there are some fleeting performance boosts: an increase of 1.8–2.1% on MGSM and mild improvements on FLORES with W8 on Aya models, and a similar translation boost of the 35B Command model at W8A8. Quantization generally has no effect or causes mild improvement on the monolingual language confusion task, and cross-lingual language confusion performance is boosted with greater quantization in some cases.
4.3 By Language
Are languages differently affected by quantization?
[…]
W4-g causes considerable degradation across languages for the 35B Command model. A relationship between language and performance is apparent: $\neg$¬Ltn/IE languages typically degrade more. Chinese, Japanese, and Korean are particularly harmed by W4 quantization of the 103B. The effect is seen consistently across all automatic metrics for Command, with limited exception. Table 5 is discussed more thoroughly in Section 4.5, but also shows this discrepancy. In the Appendix, we see the same for Aya 23 models at W4.
Interestingly, W8A8 of the 35B Command model helps on average across all languages. The magnitude is primarily due to an increase on cross-lingual language confusion. W8 also aids Aya 23 on MGSM (Table A6) for $\neg$¬Ltn/IE languages, and across languages on FLORES (Table A16).
4.4 By Model Size
How do model size and quantization level interact?
Across evaluations at the most extreme quantization (W4/W4-g), smaller models are more sensitive: W4-g variants of 103B and 35B Command record $-0.9\%$ and $-2.8\%$ performance relative to FP16 on average, with a stark difference of $-2.9\%$ vs. $-13.1\%$ on MGSM. Aya 23 35B/8B record $-2.9\%$ vs. $-3.7\%$ on average, with their largest gap occurring in Belebele ($-5.9\%$ vs. $-8.5\%$). (Refer back to Tables 1 and 2.)
4.5 By Quantization Strategy
How do techniques like SmoothQuant and group-wise scaling affect downstream performance?
FLORES Language Confusion
Avg. Rel. % mMMLU MGSM En $\rightarrow$ L2 L2 $\rightarrow$ En Monolingual Cross-lingual
Ltn/IE $\neg$ Ltn/IE $\neg$ Ltn/IE $\neg$ Ltn/IE $\neg$ Ltn/IE $\neg$ Ltn/IE $\neg$ Ltn/IE $\neg$
W8A8 -0.7% -1.0% -1.3% -2.1% -0.9% -1.3% -0.1% 0.1% -1.0% -1.6% 0.0% 0.4% -0.9% -1.6%
W8A8-sq -0.4% -0.7% -0.4% -0.8% -1.3% -1.9% 0.2% 0.0% -1.1% -1.6% -0.1% 0.1% 0.1% 0.0%
W4 -1.9% -3.3% -3.9% -4.9% -8.0% -10.2% -1.3% -2.0% -1.1% -2.3% 0.1% 0.1% 2.9% -0.4%
W4-g -0.6% -1.4% -1.1% -1.9% -1.8% -4.9% 0.2% 0.7% -0.3% -0.8% 0.1% -0.1% -0.9% -1.3%
Table 5: Effect of mitigation strategies on W8A8 and W4 quantization on the 103B model. Percentage points off FP16 baseline for W8A8-sq vs. naive W8A8 and W4-g vs. W4, broken down by Latin-script/Indo-European languages (Ltn/IE) versus others ($\neg$). Avg. Rel. % reports averaged performance all datasets.
fr es ja ko Avg Ltn/IE $\neg$
LLM RM LLM RM LLM RM LLM RM LLM RM LLM RM LLM RM
W8 1.0% -0.7% -10.2% 7.5% -5.4% 5.4% 7.5% -5.8% -1.8% 1.6% -4.6% 3.4% 1.0% -0.2%
W8A8-sq -18.4% -5.1% -3.7% 4.1% 2.0% 4.7% 3.7% -5.1% -4.1% -0.3% -11.0% -0.5% 2.9% -0.2%
W4-g -10.5% -17.0% -16.6% 2.0% -15.3% 0.0% -5.8% -15.6% -12.1% -7.7% -13.6% -7.5% -10.5% -7.8%
Internal W4 -30.2% -20.4% -33.0% -17.0% -21.7% -20.0% -18.6% -27.6% -25.9% -21.2% -31.6% -18.7% -20.2% -23.8%
W8 -1.3% 2.0% 7.3% -4.0% -6.0% -5.3% 2.7% 2.0% 0.7% -1.3% 3.0% -1.0% -1.7% -1.7%
W8A8-sq -15.3% -8.7% 8.7% -8.0% -1.3% 1.3% -8.0% -4.7% -4.0% -5.0% -3.3% -8.3% -4.7% -1.7%
W8A8 -3.4% 2.7% 13.3% -3.3% 2.7% -1.3% 5.3% -3.3% 4.5% -1.3% 5.0% -0.3% 4.0% -2.3%
Dolly W4-g -7.4% -2.7% -4.0% 4.7% -15.3% -15.3% -11.3% -5.3% -9.5% -4.7% -5.7% 1.0% -13.3% -10.3%
Table 6: Relative performance vs. FP16 of 103B quantized models according to LLM/RM-as-a-Judge over Internal and Aya Dolly subsampled test sets. Raw win-rates in Table A20.
Table 5 shows the effect of using SmoothQuant and Group-Wise scaling strategies. We evaluate variants of the 103B Command model with SmoothQuant (W8A8-sq), and a more naive W4 variant using per-column quantization instead of group-wise scaling. We compare W8A8-sq to W8A8, and W4-g to W4.
[…]
On cross-lingual language confusion, strategies aimed to retain performance have different effects: SmoothQuant recovers all lost from naive W8A8, but Group-Wise scaling is actively damaging. In contrast, W4 benefits Ltn/IE and Arabic on cross-lingual language confusion, but worsens the rest. [^11] Full results in are Table A19.
Thus, while the quantization strategies tend to aid performance overall, there may be adverse effects on specific tasks. More research is needed to understand this, but it is intriguing to consider the effect that lower-precision might have on the ability to produce output in a desired language, and maintain that language once decoding begins.
4.6 LLM/RM-as-a-Judge
Table 6 shows relative performance of quantized variants of the 103B Command model evaluated with LLM- and RM-as-a-Judge. [^12] Calculation: Quantized Win $\frac{\text{Quantized Win Rate}-50}{50}$, as 50 is the expected win-rate of two FP16 models compared. In nearly all cases, the LLM and RM agree that W4 and W4-g severely harm performance on our challenging Internal test set. Performance is also severely degraded for $\neg$¬Ltn/IE languages on Dolly with W4-g, and French with W8A8-sq. On average across languages, the LLM and RM agree on the ranking of model quality over Internal. Results on the easier Dolly test set are less clear-cut: The LLM reports greater degradation for Internal than Dolly overall, but the RM disagrees for W8 and W8A8-sq. Perhaps Dolly prompts are easy enough that models output similar responses, creating more noise in the judgments; future work could examine this hypothesis. Furthermore, on multiple instances, the LLM and RM disagree on whether performance improves or worsens, given the same setting. Comparisons between the two differing methods of automated evaluation are worthy of further study.
4.7 Human Evaluation
fr es ja ko Avg Ltn/IE $\neg$
W8 -7.4% 0.6% 7.4% -12.0% -2.8% -3.4% -2.3%
W8A8-sq -9.4% -7.4% -2.0% 4.0% -3.7% -8.4% 1.0%
Internal W4-g -16.6% -4.6% -16.0% -4.6% -10.5% -10.6% -10.3%
W8 0.6% -5.4% 12.0% 0.0% 1.8% -2.4% 6.0%
W8A8-sq -7.4% -8.6% 0.0% -3.4% -4.8% -8.0% -1.7%
Dolly W4-g -9.4% -1.4% 2.6% -8.0% -4.1% -5.4% -2.7%
Table 7: Relative performance vs. FP16 of 103B quantized models according to human evaluators over Internal and Aya Dolly subsampled test sets.
Human evaluation paints a similar picture in Table 7, with some outliers. Average performance drops steadily across evaluated languages on the Internal test set, which has more difficult prompts. The sharpest decline is in French, with $-16.6\%$ at W4-g. Curiously, there is an initial $7.4\%$ boost for Japanese with W8, but it falls to $-16.0\%$ with more extreme quantization. Interestingly, human annotators generally prefer outputs of quantized models on Dolly prompts in Japanese, too, but disprefer those in other languages. We see more pronounced degradation on Internal overall, with an average relative drop of 5.7% versus 2.4% for Dolly.
5 Related Work
Impact of Compression on Multilingual Tasks
There is a scarcity of research examining the impact of compression and quantization on multilingual tasks. Paglieri et al. (2024) study the impact of multilingual calibration sets on quantization, but their evaluation is English-only. Ramesh et al. (2023) study the effect of compression on multilingual model fairness in terms of classification accuracy, showing that while monolingual evaluation indicates a negative impact, multilingual evaluation differs across languages and dimensions. Kharazmi et al. (2023) show that recovering compression-caused performance loss of LSTMs is harder in a multilingual setting than monolingually. In machine translation, Diddee et al. (2022) show that distillation has a varied effect by language due to dependence on priors such as amount of synthetic data used and confidence of the teacher models, while quantization exhibits more consistent performance trends across languages. Our work is the first, to our knowledge, to study the effect of quantization on LLMs and for open-ended generation.
More broadly, multilingual data is an example of long tail data. Prior work shows that compression techniques like quantization and sparsity amplify disparate treatment of long-tail rare features (Hooker et al., 2019; Ahia et al., 2021; Ogueji et al., 2022; Hooker et al., 2020). Ogueji et al. (2022) show that depending on how out of distribution the task data is, sparsity-based compression can sometimes avoid overfitting to the training data, making a model better suited to the downstream task. Ahia et al. (2021) find that sparsity preserves machine translation performance on frequent sentences, but disparately impacts infrequent sentences.
Quantization of LLMs
A recent line of work has emerged on techniques to improve performance of quantized LLMs, with the sole focus on English models and data for tuning and evaluation (Ahmadian et al., 2024; Dettmers et al., 2022; Xiao et al., 2023; Bondarenko et al., 2024; Gong et al., 2024). Even the most recent (Li et al., 2024; Liu et al., 2024) omit the multilingual dimension without acknowledging the limitation. Multilinguality and compression are both integral parts of LLMs, and our work explores this new territory.
Model design choices
We consider how design choices such as quantization impact performance for users of different languages. A wider body of work examines how design choices impact performance on underrepresented features or subgroups. Zhuang et al. (2021) and Nelaturu et al. (2023) find that hardware choice incurs disparate impact on underrepresented features. Wang et al. (2022) establish that distillation imposes similar trade-offs, but the disproportionate harm to the long-tail could be mitigated by modifying the student-teacher objective. Ko et al. (2023) evaluate the positive role of ensembling disproportionately favoring underrepresented attributes. Bagdasaryan and Shmatikov (2019) show that differential privacy techniques like gradient clipping and noise injection disproportionately impact underrepresented features.
6 Conclusion & Future Work
We examine widely adopted quantization techniques for model compression and ask, How does quantization impact different languages? We perform an extensive study of quantization in state-of-the-art multilingual LLMs—from 8 billion to 103 billion parameters—in 20+ languages using automatic metrics, LLM-as-a-Judge, RM-as-a-Judge, and human evaluation. We find that: (1) Damage from quantization is much worse than appears from automatic metrics: even when not observed automatically, human evaluators notice it. (2) Quantization affects languages to varying degrees, with non-Latin script languages more severely affected on automatic benchmarks. (3) Challenging tasks degrade fast and severely: math performance is strikingly reduced, as are responses on realistic challenging prompts judged by humans. On a bright note, quantization occasionally brings performance benefits.
Our results urge attention to multilingual performance at all stages of system design. Researchers might extend our work to consider the impact of other decisions on multilingual performance, including on languages excluded from training and out-of-distribution tasks. By being mindful of the impact on long-tail features, we’ll build better systems to serve the world.
7 Limitations
Generality of findings
Due to the number of methods, languages, and benchmarks we examine, we focus our evaluation on models from two families (Command R and Aya). As we observe similar trends across these models, our findings are likely to generalize to other LLMs. Nevertheless, models that have been optimized differently or trained with a focus on specific tasks such as code or mathematical reasoning may behave differently.
Under-represented languages
For our study, we focused on languages that were supported by the models we evaluated. Performance deterioration is likely even larger for languages that are not or severely under-represented in the pre-training data. For such languages, evaluation is also more challenging due to poor availability of benchmark data and human annotators.
Appendix A Appendix
A.1 Prompts for mMMLU and LLM-as-a-Judge
The following are multiple choice questions (with answers)
about clinical knowledge.
{CJK*}UTF8mj다음 중 파제트병에 대한 설명으로 옳은 것은 무엇입니까?
{CJK*}UTF8mjA. 긴 뼈가 휘어지는 것이 특징
{CJK*}UTF8mjB. 척수압박은 흔한 합병증이다
{CJK*}UTF8mjC. 심부전은 알려진 합병증이 아니다
{CJK*}UTF8mjD. 병적 골절은 특징이 아닙니다.
Answer: B
Table A1: mMMLU prompt. Following Achiam et al. (2023), letter choices and “Answer” are kept in English.
Example Prompt
I want you to create a leaderboard of different large-language models. To do so, I will give you the conversations (prompts) given to the models, and the responses of two models. Please rank the models based on which responses would be preferred by humans. All inputs and outputs should be python dictionaries.
Here is the prompt:
       "conversation": """User: La tomate est-elle un fruit ou un légume?""",
Here are the outputs of the models:
        "model": "model_1",
        "answer": """La tomate est un fruit. Plus précisément, il s’agit d’un fruit charnu, issu de la transformation de l’ovaire de la fleur du plant de tomate."""
        "model": "model_2",
        "answer": """La tomate est un fruit du point de vue botanique, car elle contient des graines et se développe à partir de la fleur d’une plante. Cependant, en cuisine, on considère souvent la tomate comme un légume en raison de son utilisation dans des plats salés et de sa saveur moins sucrée par rapport à d’autres fruits."""
Now please rank the models by the quality of their answers, so that the model with rank 1 has the best output. Then return a list of the model names and ranks, i.e., produce the following output:
       {’model’: <model-name>, ’rank’: <model-rank>},
       {’model’: <model-name>, ’rank’: <model-rank>}
Your response must be a valid Python dictionary and should contain nothing else because we will directly execute it in Python. Please provide the ranking that the majority of humans would give.
Table A2: Example Input for LLM-as-a-Judge. Template derived from Li et al. (2023b): https://github.com/tatsu-lab/alpaca_eval/blob/main/src/alpaca_eval/evaluators_configs/gpt-3.5-turbo-1106_ranking/ranking_prompt.txt
A.2 Automatic Tasks - Full Results
de es fr ja zh Avg
103B FP16 72.6 76.6 70.6 63.0 70.2 70.6
W8 72.8 75.9 69.5 61.3 70.2 69.9
W8A8-sq 73.4 73.8 69.6 62.9 67.7 69.5
W8A8 74.1 73.9 69.8 63.4 68.0 69.8
W4-g 71.2 75.7 69.0 58.0 68.9 68.6
W4 64.6 71.3 66.5 56.1 63.5 64.4
35B FP16 56.6 57.3 51.8 38.8 44.4 49.8
W8 55.9 56.6 52.1 37.4 45.1 49.4
W8A8 54.2 53.4 49.9 35.8 42.0 47.1
W4-g 47.2 51.0 47.1 34.3 36.7 43.3
Table A3: Command model MGSM results. (Acc.)
de es fr ja zh Avg Ltn/IE $\neg$
W8 0.3% -0.9% -1.6% -2.7% -0.1% -1.0% -0.7% -1.4%
W8A8-sq 1.1% -3.7% -1.5% -0.1% -3.6% -1.6% -1.3% -1.9%
W8A8 2.1% -3.5% -1.1% 0.6% -3.2% -1.0% -0.9% -1.3%
W4-g -1.9% -1.3% -2.3% -7.9% -1.9% -3.0% -1.8% -4.9%
103B W4 -11.0% -7.0% -5.9% -10.9% -9.6% -8.8% -8.0% -10.2%
W8 -1.3% -1.1% 0.6% -3.7% 1.6% -0.8% -0.6% -1.0%
W8A8 -4.4% -6.8% -3.6% -7.6% -5.4% -5.6% -4.9% -6.5%
35B W4-g -16.7% -10.9% -9.0% -11.5% -17.3% -13.1% -12.2% -14.4%
Table A4: Relative performance (%$\Delta$) vs. FP16 for Command Models on MGSM. Ltn/IE: Latin-script/Indo-European languages: de, es, fr. $\neg$: ja, zh.
Avg (-en) Avg de en es fr ja ru zh
FP16 51.2 53.7 61.6 68.4 58.4 55.6 22.8 58.0 50.8
Aya-23-35b W8 52.1 54.2 54.4 66.4 61.2 60.4 24.4 57.2 55.2
W4 48.1 50.7 58.8 66.0 54.8 54.8 18.4 53.6 48.4
FP16 34.7 36.6 40.4 48.0 45.2 38.8 12.8 38.0 32.8
Aya-23-8b W8 35.4 36.9 39.6 45.6 45.6 38.8 13.6 38.8 36.0
W4 32.1 33.5 39.6 42.4 42.0 34.0 7.2 33.6 36.0
Table A5: Aya 23 language-specific results for MGSM (5-shot).
de en es fr ja ru zh Avg Avg (-en) Ltn/IE $\neg$
W8 -11.7% -2.9% 4.8% 8.6% 7.0% -1.4% 8.7% 1.9% 2.7% -0.3% 4.8%
Aya-23-35b W4 -4.5% -3.5% -6.2% -1.4% -19.3% -7.6% -4.7% -6.8% -7.3% -3.9% -10.5%
W8 -2.0% -5.0% 0.9% 0.0% 6.2% 2.1% 9.8% 1.7% 2.8% -1.5% 6.0%
Aya-23-8b W4 -2.0% -11.7% -7.1% -12.4% -43.8% -11.6% 9.8% -11.2% -11.2% -8.3% -15.2%
Table A6: Relative performance (%$\%\Delta$) vs. FP16 for Aya Models on MGSM (5-shot). Ltn/IE are Latin-script/Indo-European languages: de, en, es, fr. $\neg$ are the rest: ja, ru, zh.
Avg (-en) Avg ar cs de el en es fa fr hi id it ja ko nl pl pt ro ru tr uk vi zh
35b FP16 77.6 77.9 78.9 78.2 77.1 76.4 84.7 81.0 75.8 81.9 65.6 77.8 79.8 75.9 73.3 77.7 75.8 83.8 78.9 79.6 74.1 77.6 78.3 81.2
W8 77.1 77.4 77.3 78.8 77.2 76.6 84.6 80.8 74.9 82.4 65.6 77.6 80.8 74.8 73.7 77.6 74.8 82.9 77.1 78.9 72.0 77.2 77.0 80.3
W4 73.0 73.4 73.8 74.9 73.2 70.8 83.2 77.0 71.4 78.1 61.0 73.9 76.2 71.7 67.4 73.0 70.4 80.1 74.3 73.3 68.2 73.0 71.4 78.8
8b FP16 64.8 65.3 65.6 61.9 65.6 64.0 77.0 67.0 63.6 69.6 54.3 67.4 65.7 65.2 61.7 63.8 61.3 69.1 65.7 69.7 58.1 66.8 62.3 72.2
W8 64.6 65.1 64.3 61.8 64.8 63.0 76.1 67.4 63.9 70.4 54.2 67.4 64.6 65.4 61.4 64.3 59.8 68.7 65.4 68.7 58.1 67.0 63.7 71.8
W4 59.3 59.9 61.9 57.0 61.6 57.7 73.8 61.1 58.2 65.7 49.8 64.7 58.3 60.7 51.1 60.7 54.9 62.0 59.8 63.9 50.1 61.0 58.8 66.2
Table A7: Aya 23 language-specific results for Belebele. (Accuracy)
ar cs de el en es fa fr hi id it ja ko nl pl pt ro ru tr uk vi zh Avg Avg (-en) Ltn $\neg$
W8 -2.0% 0.7% 0.1% 0.2% -0.1% -0.3% -1.2% 0.7% 0.0% -0.3% 1.3% -1.5% 0.5% -0.1% -1.3% -1.1% -2.3% -0.8% -2.8% -0.4% -1.7% -1.1% -0.6% -0.6% -0.6% -0.7%
35b W4 -6.5% -4.3% -5.0% -7.4% -1.7% -4.9% -5.7% -4.6% -7.0% -5.0% -4.5% -5.6% -8.0% -6.0% -7.0% -4.4% -5.8% -7.8% -7.9% -5.9% -8.8% -3.0% -5.8% -6.0% -5.4% -6.3%
W8 -1.9% -0.2% -1.2% -1.6% -1.2% 0.7% 0.5% 1.3% -0.2% 0.0% -1.7% 0.3% -0.4% 0.9% -2.5% -0.6% -0.4% -1.4% 0.0% 0.3% 2.1% -0.6% -0.3% -0.3% -0.2% -0.5%
8b W4 -5.6% -7.9% -6.1% -9.9% -4.2% -8.8% -8.4% -5.6% -8.4% -4.1% -11.2% -7.0% -17.1% -4.9% -10.5% -10.3% -9.0% -8.3% -13.8% -8.7% -5.7% -8.3% -8.3% -8.5% -7.8% -9.1%
Table A8: Relative performance (%$\%\Delta$) vs. FP16 for Aya Models on Belebele. Ltn are Latin-script languages: cs, de, es, en, es, fr, id, it, nl, pl, pt, ro, tr, vi. $\neg$ are the rest.
ar de es fr it ja ko pt zh Avg
103B FP16 64.0 68.3 68.7 68.0 69.3 64.4 62.3 70.0 65.0 66.7
W8 64.1 68.3 68.7 68.1 69.4 64.3 62.3 69.9 65.0 66.7
W8A8-sq 63.5 67.9 68.8 68.0 69.1 63.6 61.8 69.2 64.9 66.3
W8A8 62.6 67.1 68.2 67.4 68.3 62.9 60.8 68.7 64.1 65.6
W4-g 62.9 67.5 68.2 67.6 68.6 62.8 61.1 68.6 64.0 65.7
W4 60.5 65.7 66.5 65.4 66.6 61.1 59.3 66.7 62.1 63.8
35B FP16 56.5 60.7 62.3 61.8 62.0 56.4 54.8 62.0 57.9 59.4
W8 56.5 60.6 62.2 61.8 61.9 56.4 54.7 62.1 57.9 59.3
W8A8 56.4 60.5 62.5 61.9 62.0 55.8 54.5 61.8 58.1 59.3
W4-g 55.4 59.7 62.0 61.0 60.7 54.4 53.2 60.8 56.6 58.2
Table A9: mMMLU scores for Command Models. (Accuracy)
ar de es fr it ja ko pt zh Avg Ltn/IE $\neg$
W8 0.2% 0.0% 0.0% 0.1% 0.1% -0.2% 0.0% -0.1% 0.0% 0.0% 0.0% 0.0%
W8A8-sq -0.8% -0.6% 0.1% 0.1% -0.3% -1.3% -0.8% -1.1% -0.2% -0.5% -0.4% -0.8%
W8A8 -2.2% -1.8% -0.7% -1.0% -1.5% -2.3% -2.4% -1.8% -1.4% -1.7% -1.3% -2.1%
W4-g -1.7% -1.2% -0.7% -0.6% -1.0% -2.5% -1.9% -2.0% -1.5% -1.5% -1.1% -1.9%
103B W4 -5.5% -3.8% -3.1% -3.9% -3.8% -5.1% -4.8% -4.8% -4.4% -4.4% -3.9% -4.9%
W8 0.0% -0.2% -0.2% 0.0% -0.2% 0.0% -0.2% 0.2% 0.0% -0.1% -0.1% 0.0%
W8A8 -0.2% -0.3% 0.3% 0.2% 0.0% -1.1% -0.5% -0.3% 0.3% -0.2% 0.0% -0.4%
35B W4-g -1.9% -1.6% -0.5% -1.3% -2.1% -3.5% -2.9% -1.9% -2.2% -2.0% -1.5% -2.7%
Table A10: Relative performance (%$\Delta$) vs. FP16 for Command Models on mMMLU. Ltn/IE are Latin-script/Indo-European languages: de, es, fr, it, pt. $\neg$ are the rest: ar, ja, ko, zh.
Avg (-en) Avg ar de en es fr hi id it nl pt ro ru uk vi zh
FP16 58.2 58.8 53.9 60.4 66.7 61.6 62.0 47.8 58.9 61.5 60.3 62.0 59.7 57.8 56.3 55.3 57.5
Aya-23-35b W8 57.9 58.5 53.8 60.0 66.2 61.7 61.7 47.4 58.7 61.1 60.0 61.6 59.1 57.5 56.1 54.9 57.5
W4 56.6 57.2 52.3 58.7 65.2 60.3 60.4 45.7 57.4 59.8 58.6 60.5 57.7 56.5 55.0 53.8 56.1
FP16 48.2 48.6 45.1 50.0 54.6 50.9 51.0 39.7 48.8 50.7 49.7 50.8 49.9 47.8 46.8 46.5 47.1
Aya-23-8b W8 47.8 48.2 44.9 49.9 54.2 50.5 50.6 39.4 48.5 50.2 49.4 50.6 49.2 47.4 46.3 45.7 46.4
W4 46.7 47.1 43.9 48.4 53.4 49.4 49.0 38.4 47.5 49.1 47.9 49.1 48.0 46.2 45.6 44.9 46.1
Table A11: Aya 23 language-specific results for mMMLU (Okapi). (Accuracy)
ar de en es fr hi id it nl pt ro ru uk vi zh Avg Avg (-en) Ltn $\neg$
W8 -0.2% -0.7% -0.6% 0.2% -0.4% -0.7% -0.3% -0.6% -0.5% -0.6% -1.1% -0.6% -0.3% -0.8% -0.1% -0.5% -0.5% -0.5% -0.4%
Aya-23-35b W4 -3.1% -2.8% -2.2% -2.0% -2.5% -4.4% -2.5% -2.7% -2.8% -2.5% -3.3% -2.2% -2.2% -2.7% -2.6% -2.7% -2.7% -2.6% -2.9%
W8 -0.5% -0.3% -0.8% -0.8% -0.8% -0.9% -0.6% -1.0% -0.7% -0.4% -1.4% -0.9% -1.1% -1.7% -1.5% -0.9% -0.9% -0.8% -1.0%
Aya-23-8b W4 -2.7% -3.3% -2.1% -2.9% -3.9% -3.4% -2.6% -3.2% -3.7% -3.3% -3.8% -3.4% -2.7% -3.5% -2.2% -3.1% -3.2% -3.2% -2.9%
Table A12: Relative performance (%$\%\Delta$) vs. FP16 for Aya Models on mMMLU (Okapi). Ltn are Latin-script languages: de, en, es, fr, id, it, nl, pt, ro, vi. $\neg$ are the rest.
English $\rightarrow$ L2 L2 $\rightarrow$ English
ar de es fr it ja ko pt zh Avg ar de es fr it ja ko pt zh Avg
103B FP16 27.1 40.0 30.1 50.6 33.1 33.1 29.1 51.0 45.1 37.7 45.0 46.3 33.4 48.6 36.5 29.5 33.0 52.2 32.1 39.6
W8 27.2 40.0 30.0 50.7 33.1 33.2 29.1 50.9 45.1 37.7 45.2 46.3 33.4 48.5 36.5 29.5 33.0 52.1 32.0 39.6
W8A8-sq 26.8 40.3 30.0 51.0 33.0 33.1 29.3 51.2 45.1 37.8 44.5 46.2 32.9 48.1 35.9 29.3 32.5 51.6 31.2 39.1
W8A8 26.9 39.8 30.0 50.9 33.0 33.7 29.0 51.1 45.1 37.7 44.4 45.9 33.1 47.9 36.2 29.2 32.5 51.8 31.4 39.1
W4-g 27.3 40.4 30.1 51.0 33.0 33.9 29.3 50.9 44.7 37.8 44.9 46.4 33.2 48.4 36.3 29.3 32.7 52.0 31.6 39.4
W4 26.9 39.1 29.9 50.0 32.8 32.8 27.9 50.3 44.0 37.1 44.2 45.8 33.1 47.9 36.0 29.0 32.3 51.8 30.9 39.0
35B FP16 20.1 33.5 27.8 44.5 29.7 27.0 22.7 45.5 40.4 32.4 38.4 41.2 31.8 43.1 34.0 26.2 28.4 48.1 28.4 35.5
W8 20.0 33.4 27.8 44.5 29.7 26.9 22.9 45.3 40.3 32.3 38.3 41.1 31.7 43.0 34.0 26.4 28.2 48.0 28.2 35.4
W8A8 21.2 34.1 27.8 45.1 30.0 27.6 23.1 46.1 40.8 32.9 38.5 42.2 31.7 43.5 34.2 26.5 28.6 48.6 28.7 35.8
W4-g 18.8 32.9 27.7 43.9 29.6 26.0 22.1 45.1 39.7 31.7 38.3 41.4 31.0 43.1 34.0 25.5 28.1 48.0 28.0 35.3
Table A13: Full results on FLORES for Command Models. (SacreBLEU)
English $\rightarrow$ L2 L2 $\rightarrow$ English
ar de es fr it ja ko pt zh Avg Ltn/IE $\neg$ ar de es fr it ja ko pt zh Avg Ltn/IE $\neg$
W8 0.1% 0.1% -0.4% 0.2% 0.1% 0.3% -0.2% -0.3% 0.1% 0.0% -0.1% 0.1% 0.4% -0.1% -0.1% -0.1% -0.1% -0.1% 0.0% -0.1% -0.1% 0.0% -0.1% 0.1%
W8A8-sq -1.1% 0.8% -0.4% 0.7% -0.2% 0.2% 0.7% 0.3% 0.1% 0.1% 0.2% 0.0% -1.2% -0.1% -1.4% -1.0% -1.8% -0.7% -1.6% -1.1% -2.9% -1.3% -1.1% -1.6%
W8A8 -1.0% -0.4% -0.5% 0.5% -0.4% 1.8% -0.4% 0.1% 0.0% 0.0% -0.1% 0.1% -1.3% -0.8% -1.1% -1.4% -1.0% -1.2% -1.7% -0.8% -2.1% -1.3% -1.0% -1.6%
W4-g 0.7% 1.0% -0.3% 0.8% -0.3% 2.6% 0.5% -0.3% -0.8% 0.4% 0.2% 0.7% -0.3% 0.2% -0.6% -0.3% -0.6% -0.7% -0.9% -0.3% -1.4% -0.6% -0.3% -0.8%
103B W4 -0.8% -2.2% -0.7% -1.3% -0.9% -0.8% -4.3% -1.5% -2.3% -1.6% -1.3% -2.0% -1.8% -1.1% -0.8% -1.4% -1.6% -1.7% -2.2% -0.7% -3.6% -1.7% -1.1% -2.3%
W8 -0.7% -0.4% -0.1% 0.0% 0.0% -0.2% 0.7% -0.4% -0.2% -0.1% -0.2% -0.1% -0.2% -0.2% -0.5% -0.3% 0.0% 0.8% -0.5% -0.1% -0.6% -0.2% -0.2% -0.1%
W8A8 5.5% 1.9% 0.1% 1.4% 0.9% 2.1% 1.9% 1.4% 0.9% 1.8% 1.1% 2.6% 0.5% 2.5% -0.6% 0.9% 0.5% 1.1% 0.8% 1.1% 1.0% 0.9% 0.9% 0.8%
35B W4-g -6.7% -1.9% -0.4% -1.3% -0.4% -3.9% -2.8% -0.7% -1.7% -2.2% -1.0% -3.8% -0.1% 0.6% -2.5% 0.0% -0.1% -2.8% -1.1% -0.2% -1.4% -0.8% -0.5% -1.3%
Table A14: Relative performance (%$\Delta$) vs. FP16 for Command Models on FLORES. Ltn/IE are Latin-script/Indo-European languages: de, es, fr, it, pt. $\neg$ are the rest: ar, ja, ko, zh.
[…]
ar cs de el es fa fr he hi id it ja ko nl pl pt ro ru tr uk vi zh Avg
FP16 40.0 39.1 42.5 36.3 32.1 33.4 54.1 39.5 31.9 44.7 36.6 28.7 25.5 33.4 30.7 53.1 43.3 38.9 33.8 38.2 41.0 34.0 37.8
Aya-23-35b W8 40.0 39.0 42.9 36.2 32.2 33.7 53.9 40.0 32.3 44.8 36.5 28.9 25.5 33.6 30.9 53.2 43.4 38.7 33.8 38.3 41.4 33.8 37.9
W4 39.3 38.0 42.5 36.0 32.0 32.6 53.3 39.1 31.2 44.6 36.1 28.2 25.1 32.9 30.0 52.8 42.6 38.3 33.2 37.8 40.8 33.1 37.3
FP16 36.3 35.7 39.3 34.0 31.5 30.0 51.0 35.0 27.2 43.4 34.7 24.9 22.0 32.2 28.4 50.2 41.6 35.0 29.1 34.2 39.0 30.1 34.8
Aya-23-8b W8 36.5 36.1 39.5 33.9 31.4 30.4 51.4 35.0 27.0 43.2 34.8 24.8 22.2 32.1 28.5 50.0 42.0 34.9 28.9 34.3 39.0 30.6 34.8
W4 35.4 35.0 39.2 33.4 31.2 29.6 50.2 33.3 26.4 42.8 34.3 24.3 21.5 31.8 28.0 49.8 40.9 34.2 28.1 33.7 38.5 29.4 34.1
[…]
FP16 46.4 45.3 48.9 42.4 37.7 41.3 50.6 48.3 42.7 48.5 40.5 33.7 35.3 37.7 36.4 54.8 49.5 41.6 42.2 44.8 41.4 34.8 42.9
Aya-23-35b W8 46.4 45.4 49.0 42.2 37.3 41.4 50.7 48.6 42.9 48.7 40.5 34.0 35.1 37.5 36.4 54.8 49.5 41.7 42.2 45.0 41.5 34.9 43.0
W4 45.7 44.9 48.5 41.8 37.5 40.5 50.4 47.3 41.8 48.0 40.8 33.1 34.4 37.2 35.8 54.3 49.2 41.6 41.4 44.2 40.9 34.2 42.4
FP16 42.4 42.0 46.5 38.7 35.4 36.5 48.1 43.7 37.4 45.5 37.9 29.9 30.9 35.8 33.6 51.7 46.7 38.6 36.9 41.2 38.2 31.6 39.5
Aya-23-8b W8 42.1 42.5 46.7 39.2 35.5 36.8 48.1 44.2 37.7 45.5 38.2 30.0 31.3 35.6 33.7 52.0 46.6 38.5 37.0 41.6 38.4 31.9 39.7
W4 41.4 42.2 46.2 38.1 35.8 36.7 47.4 42.8 36.2 44.7 38.5 29.7 30.1 35.7 33.1 51.9 46.1 38.3 35.7 40.7 37.6 31.6 39.1
Table A15: Aya 23 language-specific results for FLORES. spBLEU with FLORES200 tokenizer.
[…]
ar cs de el es fa fr he hi id it ja ko nl pl pt ro ru tr uk vi zh Avg Ltn $\neg$
W8 -0.1% -0.2% 1.1% -0.1% 0.2% 0.7% -0.4% 1.2% 1.3% 0.3% -0.4% 0.6% 0.0% 0.6% 0.7% 0.3% 0.4% -0.6% 0.1% 0.2% 0.9% -0.8% 0.3% 0.3% 0.3%
Aya-23-35b W4 -1.7% -2.7% 0.0% -0.7% -0.4% -2.4% -1.4% -1.0% -2.0% -0.2% -1.5% -2.0% -1.5% -1.4% -2.2% -0.5% -1.6% -1.6% -1.8% -1.1% -0.5% -2.7% -1.4% -1.2% -1.7%
W8 0.5% 1.1% 0.6% -0.3% -0.4% 1.3% 0.8% -0.1% -0.9% -0.3% 0.2% -0.2% 1.0% -0.1% 0.2% -0.5% 0.9% -0.1% -0.7% 0.0% 0.0% 1.7% 0.2% 0.2% 0.3%
Aya-23-8b W4 -2.5% -2.0% -0.3% -1.7% -1.0% -1.4% -1.6% -5.1% -3.1% -1.2% -1.3% -2.1% -2.0% -1.1% -1.3% -0.9% -1.8% -2.1% -3.6% -1.7% -1.1% -2.2% -1.9% -1.4% -2.4%
[…]
W8 0.0% 0.3% 0.2% -0.5% -1.1% 0.3% 0.3% 0.7% 0.4% 0.3% 0.0% 0.8% -0.7% -0.5% 0.1% -0.1% -0.1% 0.3% -0.1% 0.4% 0.2% 0.2% 0.1% 0.0% 0.2%
Aya-23-35b W4 -1.6% -0.9% -1.0% -1.4% -0.4% -2.0% -0.4% -2.0% -2.2% -1.1% 0.8% -1.6% -2.7% -1.5% -1.5% -0.9% -0.5% 0.1% -2.0% -1.5% -1.4% -2.0% -1.2% -0.9% -1.7%
W8 -0.6% 1.3% 0.4% 1.3% 0.3% 0.8% 0.1% 1.1% 0.6% 0.0% 0.7% 0.3% 1.4% -0.4% 0.4% 0.6% -0.1% -0.3% 0.3% 0.8% 0.6% 0.8% 0.5% 0.4% 0.6%
Aya-23-8b W4 -2.2% 0.6% -0.6% -1.6% 1.2% 0.4% -1.3% -2.1% -3.2% -1.8% 1.5% -0.6% -2.7% -0.3% -1.3% 0.5% -1.2% -0.9% -3.2% -1.4% -1.6% -0.1% -1.0% -0.6% -1.4%
Table A16: Relative performance (%$\%\Delta$) vs. FP16 for Aya Models on FLORES. Ltn are Latin-script languages: cs, de, es, fr, id, it, nl, pl, pt, ro, tr, vi. $\neg$ are the rest.
Avg XSC XCOPA XWNG
FP16 70.8 65.1 62.8 84.4
Aya-23-35b W8 70.6 65.0 62.9 83.9
W4 70.5 64.8 62.3 84.5
FP16 67.6 62.3 59.8 80.7
Aya-23-8b W8 67.6 62.4 60.0 80.6
W4 67.5 62.3 59.6 80.6
Table A17: Performance of quantized Aya 23 models on unseen discriminative tasks. XStoryCloze (XSC), XCOPA, and XWinograd (XWNG).
Monolingual Cross-Lingual
ar de es fr it ja ko pt zh Avg ar de es fr it ja ko pt zh Avg
103B FP16 99.3 100.0 99.3 99.6 100.0 98.6 100.0 98.3 97.9 99.2 93.0 90.6 91.2 91.6 93.0 93.1 91.1 88.3 91.3 91.5
W8 99.0 100.0 99.5 99.4 99.8 99.2 99.8 97.8 98.5 99.2 92.6 91.1 91.7 91.4 92.9 92.8 91.3 87.4 89.7 91.2
W8A8-sq 99.4 100.0 99.3 99.6 100.0 98.6 100.0 97.7 98.4 99.2 93.3 91.5 91.4 92.4 92.1 93.3 92.1 87.6 90.0 91.5
W8A8 99.3 100.0 99.5 99.8 100.0 99.0 99.8 98.1 99.1 99.4 91.3 89.3 91.0 91.1 91.8 93.0 89.3 87.3 89.2 90.4
W4-g 99.1 100.0 99.6 99.9 100.0 97.4 100.0 98.1 98.9 99.2 90.6 89.9 90.7 91.7 93.1 92.8 90.6 85.4 89.6 90.5
W4 99.4 100.0 99.4 99.7 99.8 99.6 99.0 98.9 98.4 99.3 95.8 94.3 95.9 93.8 93.6 92.6 88.9 90.5 89.7 92.8
35B FP16 99.2 97.0 98.1 99.2 99.6 99.6 99.0 99.0 97.7 98.7 58.8 59.6 69.0 73.0 63.6 66.3 69.2 64.2 74.6 66.5
W8 99.7 97.0 98.1 98.9 100.0 99.8 99.0 99.3 97.4 98.8 59.8 58.6 69.2 72.8 62.3 66.8 68.7 64.3 74.5 66.3
W8A8 99.9 98.0 97.1 98.9 100.0 100.0 100.0 99.0 98.4 99.0 61.0 63.9 72.1 75.1 66.2 68.3 70.1 67.4 76.3 68.9
W4-g 99.4 95.0 96.5 99.9 100.0 99.8 97.0 98.3 98.6 98.3 60.4 59.3 72.8 73.3 64.8 65.4 70.4 64.6 73.0 67.1
Table A18: Language Confusion scores for Command Models. (Line-level pass rate (LPR))
Monolingual
ar de es fr it ja ko pt zh Avg Ltn/IE $\neg$
103B W8 -0.3% 0.0% 0.2% -0.2% -0.2% 0.6% -0.2% -0.5% 0.6% 0.0% -0.1% 0.2%
W8A8-sq 0.1% 0.0% 0.0% 0.0% 0.0% 0.0% 0.0% -0.6% 0.5% 0.0% -0.1% 0.1%
W8A8 0.0% 0.0% 0.2% 0.2% 0.0% 0.4% -0.2% -0.2% 1.2% 0.2% 0.0% 0.4%
W4-g -0.2% 0.0% 0.3% 0.4% 0.0% -1.2% 0.0% -0.2% 1.0% 0.0% 0.1% -0.1%
W4 0.0% 0.0% 0.1% 0.1% -0.2% 1.0% -1.0% 0.6% 0.5% 0.1% 0.1% 0.1%
35B W8 0.5% 0.0% 0.0% -0.3% 0.4% 0.2% 0.0% 0.3% -0.3% 0.1% 0.1% 0.1%
W8A8 0.7% 1.0% -0.9% -0.3% 0.4% 0.4% 1.0% 0.0% 0.7% 0.3% 0.0% 0.7%
W4-g 0.2% -2.1% -1.6% 0.7% 0.4% 0.2% -2.0% -0.7% 0.9% -0.4% -0.6% -0.2%
Cross-lingual
103B W8 -0.5% 0.5% 0.5% -0.2% -0.1% -0.3% 0.3% -1.0% -1.8% -0.3% 0.0% -0.6%
W8A8-sq 0.3% 1.0% 0.2% 0.9% -0.9% 0.2% 1.1% -0.8% -1.4% 0.1% 0.1% 0.0%
W8A8 -1.8% -1.5% -0.2% -0.6% -1.2% -0.2% -1.9% -1.1% -2.3% -1.2% -0.9% -1.6%
W4-g -2.6% -0.8% -0.6% 0.1% 0.1% -0.3% -0.5% -3.3% -1.8% -1.1% -0.9% -1.3%
W4 3.0% 4.0% 5.1% 2.3% 0.6% -0.5% -2.4% 2.5% -1.8% 1.4% 2.9% -0.4%
35B W8 1.7% -1.6% 0.2% -0.3% -2.0% 0.8% -0.8% 0.3% -0.2% -0.2% -0.7% 0.4%
W8A8 3.8% 7.2% 4.4% 2.9% 4.1% 3.1% 1.2% 5.0% 2.3% 3.8% 4.7% 2.6%
W4-g 2.8% -0.5% 5.4% 0.5% 1.9% -1.3% 1.7% 0.6% -2.2% 1.0% 1.6% 0.3%
Table A19: Relative performance (%$\Delta$) vs. FP16 for Command Models on Language Confusion metrics. Ltn/IE are Latin-script/Indo-European languages: de, es, fr, it, pt. $\neg$ are the rest: ar, ja, ko, zh.
A.3 RM/LLM-as-a-Judge and Human Evaluation - Full Results
fr es ja ko
LLM RM LLM RM LLM RM LLM RM
W8 50.5 49.7 44.9 53.7 47.3 52.7 53.7 47.1
W8A8-sq 40.8 47.5 48.1 52.0 51.0 52.4 51.9 47.5
W4-g 44.8 41.5 41.7 51.0 42.4 50.0 47.1 42.2
Internal W4 34.9 39.8 33.5 41.5 39.2 40.0 40.7 36.2
W8 49.3 51.0 53.7 48.0 47.0 47.3 51.3 51.0
W8A8-sq 42.3 45.7 54.3 46.0 49.3 50.7 46.0 47.7
W8A8 48.3 51.3 56.7 48.3 51.3 49.3 52.7 48.3
Dolly W4-g 46.3 48.7 48.0 52.3 42.3 42.3 44.3 47.3
Table A20: LLM/RM-as-a-Judge Raw win-rates of 103B quantized models vs. FP16 over Internal and Aya Dolly subsampled test sets.
fr es ja ko
W8 46.3 50.3 53.7 44.0
W8A8-sq 45.3 46.3 49.0 52.0
Internal W4-g 41.7 47.7 42.0 47.7
W8 50.3 47.3 56.0 50.0
W8A8-sq 46.3 45.7 50.0 48.3
Dolly W4-g 45.3 49.3 51.3 46.0
Table A21: Human evaluation raw win-rates of 103B quantized models vs. FP16 over Internal and Aya Dolly subsampled test sets.